OpenStudy (anonymous):

6x - 6y = 42 7x + 6y = -68 Find the solution of this system of equations.

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@babygirl11271963

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

use elimination for this one

OpenStudy (anonymous):

its actually set up for this already

OpenStudy (anonymous):

add the 2 equations together

OpenStudy (anonymous):

and the y value will dissapear

OpenStudy (the_fizicx99):

6x - 6y = 42 7x + 6y = -68 The y's cancel out so your left with 13x = -26 /13 /13 x = -2 Plugging that into the first equation will give you, 6(-2) -6y = 42 -12 - 6y = 42 +12 +12 -6y = 54 /-6 /-6 y = -9

OpenStudy (anonymous):

Thanks @tHe_FiZiCx99 always a hero! :)

OpenStudy (anonymous):

i guess im too slow XD

OpenStudy (the_fizicx99):

6(-2) -6(-9) = 42 -12 + 54 = 42 42 = 42 therefore this is true :3

OpenStudy (the_fizicx99):

Sanku @Dakotafox79 :3